OpenStudy (anonymous):

A survey was conducted among 200 students of age groups 7-12 years and 13-18 years to find their favorite movie genre. The students had to select any one genre out of comedy, action, and adventure. Out of the 100 students in the age group 7-12 years who participated in the survey, 55 liked action or adventure movies. The total number of students of both age groups who liked comedy was 98. Using a two-way table, compute the total number of students in the age group 13-18 years who liked action or adventure movies. 43 47 53 57

OpenStudy (anonymous):

Ok so the answer is 53... you wanna know why?

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

ok so The total number of students of both age groups who liked comedy was 98. so 98/2= 49 And it says already that the 7-12 age group has 55 liked action or adventure. And 55+49=104 But there's 100 people for each age group so 104-100=4 and 49+4=53 answer:53

OpenStudy (anonymous):

if you need more help fan me or tag me :)